The weights of steers in a herd are distributed normally. The standard deviation is 200lbs and the mean steer weight is 1000lbs. Find the probability that the weight of a randomly selected steer is greater than 839lbs. Round your answer to four decimal places.
Answer #1

Given:

\mu = 1000 lbs, \sigma = 200 lbs

X : weights of steers in a herd

Find: P(X > 839 lbs)

P( X > 839) = P\left ( Z > \frac{839 - \mu }{\sigma} \right)

P( X > 839) = P\left ( Z > \frac{839 - 1000}{200} \right)

P(X > 839) = P(Z > -0.81)

P(X > 839) = 1 - P(Z < -0.81)

P(X > 839) = 1 - 0.2104 ....................From standard Normal Table

P(X > 839) = 0.7896

0.7896 is the probability that the weight of a randomly selected steer is greater than 839 lbs.

P(X> 839) P(Z> -0.81) 0.7896 z =-0.81